# Is animal magnetoreception quantum? A perspective from the energy resolution limit

### v1

@@ I Introduction
Although numerous behavioural experiments have established the ability of many species to sense the geomagnetic field, the mechanisms behind biological magnetic sensing, and their working at the physical, chemical and physiological level, are still being debated. There is evidence that migratory birds [11, 12, 13, 14, 15], fruit flies [16, 17, 18, 19, 20, 21, 22, 23, 24], butterflies [25, 26, 27], turtles [28, 30, 29], sharks [31, 32, 33], eels [34, 35, 36, 37, 38], salmon [39, 40, 41], are among the many species having magnetic sensing capabilities, either at the level of a compass for acquiring directional information towards long-distance navigation, or in some form of a magnetic map for positioning, or both. Three mechanisms have so far been prevalent in the discussion of biological magnetoreception, the radical-pair [42, 43, 44, 45, 46], the magnetite [47, 48, 49, 50, 51], and the induction mechanism [52, 53, 54, 55], while a mechanism synthesizing the first two, the MagR protein complex, was relatively recently proposed [56].
@@
Figure 1: Magnetic sensing energy resolution limit [85]. Magnetic sensitivity versus sensor’s linear dimension for tens of magnetic sensing realizations of several magnetometers, like atomic magnetometers, SQUIDs, diamond and Hall effects sensors, and more. The red solid line is defined by the energy resolution limit, that is, the bound $(\widetilde{\delta B})^{2}V/2\mu_{0}\geq\hbar$, where $\widetilde{\delta B}$ is the magnetic sensitivity (or equivalently magnetic noise) in units of ${\rm T/\sqrt{Hz}}$, $V$ is the volume of the sensor (the third power of the linear dimension), and $\mu_{0}$ vacuum’s magnetic permeability constant. If the measurement time is $\tau$, and the error in the magnetic field estimate obtained within the measurement time is $\delta B$, then $\widetilde{\delta B}=\delta B\sqrt{\tau}$. Figure reproduced from [85] with permission of the American Physical Society.
@@
Qualitatively, the closer to $\hbar$ the numerical value of the ER is for any given sensor, the more “quantum” can the sensor be characterized. This offers an alternative perspective in the study of quantum biology of magnetoreception [87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. This discussion is not about a choice between “black” or “white”, since “quantumness” can now be quantified, and spans a continuum. Indeed, there are now quantifiers for quantum coherence [97] and entanglement [98], and some have been applied to quantifying “quantumness” of biological magnetoreception [99, 100]. The ER provides yet another quantifier, this time in units of $\hbar$. For example, if it is found that some biological magentic sensor has ER e.g. $10^{7}\hbar$, and thus is far from the quantum limit, this finding is still quite informative, since it implies room for $10^{7}$ improvement. Even if Nature did not have the chance to realize such an improvement, it could in principle be engineered with biomimetic sensors [101, 102, 103].
Quantitatively, the ERL provides a strong and subtle constraint of basic and measurable parameters ($\delta B$, $V$, $\tau$). This constraint is particularly helpful for understanding biological magnetoreceptors, which unavoidably face the complexities and uncertainties of biological systems, as opposed to precisely controlled quantum magnetometers in the laboratory. As will be outlined in the paper, some biological ERs do not have a specific value as man-made sensors do, but rather, a range of possible values, stemming from those uncertainties. Conversely, it will be seen that the ERL narrows down those uncertainties. For example, we here obtain a precise lower bound for the number of cryptochromes that are required for obtaining a given magnetic sensitivity within a given measurement time.
The outline of this paper is the following. In Sec. II we make some introductory comments about the energy resolution limit aimed at an interdisciplinary readership. We then apply the ERL to the radical-pair (Sec. III), the magnetite (Sec. IV) and the MagR mechanism (Sec. V). In Sec. VI we present the results in a global ERL map. We conclude in Sec. VII.
@@ II Introductory comments about the energy resolution limit
The left-hand-side defines the energy resolution of the sensor, and the right-hand side is Planck’s constant $\hbar=10^{-34}~{}{\rm Js}$, which constant pervades quantum phenomena.
In the expression of the energy resolution, $\delta B$ is the uncertainty in the sensor’s magnetic field estimate, $V$ the sensitive volume, and $\tau$ the measurement time. The constant $\mu_{0}=4\pi\times 10^{-7}~{}{\rm Tm/A}$ is the magnetic permeability of vacuum. The name “energy resolution” derives from the following. From classical electromagnetism we know that $B^{2}/2\mu_{0}$ is the magnetic energy density (unit ${\rm J/m^{3}}$) due to the presence of a magnetic field $B$ in space. For a homogeneous magnetic field, the total energy within some volume $V$ is then $B^{2}V/2\mu_{0}$. The ERL involves the variance $(\delta B)^{2}$ instead of $B^{2}$, therefore one can loosely think of it as the minimum magnetic energy measurable within the sensor volume during the time available for the measurement. The right-hand-side of (1) is approximately equal to $\hbar$, thus it is tempting to guess that some basic quantum principle underlies the ERL.
@@
In Fig. 1 we reproduce the results of [85]. The $y$-axis displays the quantity $\widetilde{\delta B}=\delta B\sqrt{\tau}$ (having unit ${\rm T/\sqrt{Hz}}$) versus the linear dimension $L$ of the sensor. This is connected to the volume entering (1) by $V=L^{3}$. The spectral magnetic sensitivity $\widetilde{\delta B}$ describes how the measurement noise is distributed over the measurement bandwidth, given by $1/\tau$. The utility of $\widetilde{\delta B}$ is that we can produce a common figure for $\widetilde{\delta B}$ versus $L$ for a number of sensors having different measurement times $\tau$. It is seen in Fig. 1 that all sensors examined have $\widetilde{\delta B}\geq\sqrt{2\mu_{0}\hbar/L^{3}}$, which is another way to express the bound (1).
In the next three sections we analyze the energy resolution for the radical-pair, the magnetite, and the MagR mechanism. In the beginning of each section we briefly recapitulate the workings of the relevant mechanism. To establish the ER, we extract from the literature the magnetic sensitivity $\delta B$, the sensor volume $V$, and the measurement time $\tau$. At the end, we produce a similar map of the global results like the one in Fig. 1.
@@ III Radical-pair Mechanism
Within the perspective of the ERL we focus on the magnetic field amplitude. This does not limit the utility of our results, since as stated in [6], “map and compass cues might not always be as separable as previously thought”. Information on the magnetic field change is reflected in the change in concentration of radical-pair reaction products. We remind the reader that such reactions involve a coherent spin motion, which is influenced by the magnetic field. This spin motion is driven by the intra-molecular magnetic fields produced by the hyperfine couplings of the two unpaired electrons of the donor and acceptor molecule forming the radical-pair with the magnetic nuclei of the respective molecule. The radical-pair is created by photo-excitation of the neutral precursor. The mechanism is depicted in Fig. 2a, which shows the photo-excitation step of the donor-acceptor molecular dyad, DA, the electron transfer creating the radical-pair state ${\rm D^{\bullet+}A^{\bullet-}}$ in the singlet state of the two unpaired electrons, the coherent spin motion ${\rm{}^{S}D^{\bullet+}A^{\bullet-}}\leftrightarrow{\rm{}^{T}D^{\bullet+}A^{\bullet-}}$ interchanging singlet with triplet radical-pair states, and the inverse electron transfer leading to the singlet (DA) and triplet (${\rm{}^{T}DA}$) reaction products.
@@
The signalling mechanism conveying the change in reaction yields to higher levels of neurophysiological processing is still not understood [126, 100]. Therefore, as magnetic-dependent observable one usually considers one of the two reaction yields, e.g the singlet. This is found by integrating the singlet probability of the radical-pair state times the probability for a singlet recombination, which for a time interval $dt$ is given by $k_{\rm S}dt$. Similarly, the probability for triplet recombination is $k_{\rm T}dt$. The rates $k_{\rm S}$ and $k_{\rm T}$ are characteristic parameters of the specific radical-pair considered. In general they are different, however, in many simulations they are taken equal, $k_{S}=k_{T}\equiv k$. In this case, also considered herein, $\tau=1/k$ is the reaction time, which will also serve as the measurement time in our ER considerations.
@@
For completeness, in Fig. 2b we depict an example of the coherent spin motion ${\rm{}^{S}D^{\bullet+}A^{\bullet-}}\leftrightarrow{\rm{}^{T}D^{\bullet+}A^{\bullet-}}$ for an infinite radical-pair lifetime, i.e. $k=0$. We plot the singlet probability as a function of time for two different magnetic fields, to show how the field modulates the singlet-triplet mixing. In Fig. 2c we “turn on” the recombination channels, so in this case we observe the decay of the singlet probability, since radical-pairs recombine and their population tends to zero, while the population of the reaction products increases. This is shown in Fig. 2d for the singlet reaction yield, which is seen to convey the information on the magnetic field. Indeed, for a change from $B=0$ to $B=0.5~{}{\rm G}$, the singlet reaction yield at the end of the reaction ($t\gg\tau$) changes by 10% (in absolute terms).
@@ III.1 Energy resolution of synthesized radical-pairs probed with laser spectroscopy
The authors in [127] synthesized a radical-pair model using a carotenoid-porphyrin-fullerene triad (CPF). The porphyrin serves as an intermediate cation in the electron-transfer process, whereas the actual radical-pair is ${\rm C^{\bullet+}PF^{\bullet-}}$, i.e. the carotenoid is the radical cation, and the fullerene is the radical anion. The authors observed a clear magnetic field effect in the radical-pair reaction at earth’s magnetic field, albeit at 113 K, i.e. much cooler than the physiological animal temperature. Nevertheless, these data are perfectly suitable to apply the ERL. The authors measured the transient absorption of the radical-pair state at $39~{}{\rm\mu T}$ and $49~{}{\rm\mu T}$. From the noise in their data (Fig. 2c of [127]) one can estimate the magnetic sensitivity to be $\delta B\approx 2.5~{}{\rm\mu T}$. The measurement time was $\tau\approx 1~{}{\rm\mu s}$, given by the duration of the magnetic-field effect. Finally, the sensing volume of the $100~{}{\rm\mu M}$ CPF solution that was probed by the laser measuring the transient absorption of ${\rm C^{\bullet+}PF^{\bullet-}}$ was $V\approx 2\times 10^{-8}~{}{\rm m^{-3}}$ [128]. Thus in this case, it is ${\rm ER}=(\delta B)^{2}V\tau/2\mu_{0}=5\times 10^{14}\hbar$.
@@
The main reason for the large value of this ER is the large value of $\delta B$ deriving from the specific measurement scheme. It is known from quantum magnetometers working with atomic vapors [129] that the number of atoms participating in the measurement is crucial. The number density of CPF molecules, given their $100~{}{\rm\mu M}$ concentration, is $6\times 10^{16}~{}{\rm cm}^{-3}$, two orders of magnitude higher than the atom number density of typical optical pumping magnetometers [129], which deliver magnetic sensitivity at the level of 1 fT. Thus, $\delta B$ is not limited by the number of CPF molecules. Similarly, atomic magnetometers have about 3-4 orders of magnitude longer relaxation time, so again, $\tau$ is not the limiting factor here. Instead, it appears that the ER is dominated by the limited sensitivity of the specific measurement scheme.
@@
Indeed, the magnetic field effect is measured by detecting an optical absorption exciting the radical-pair state into higher electronically excited states, without any discrimination between singlet or triplet radical-pairs. That is, the magnetic field effect probes the total population of the radical-pair state, given by $\rm Tr\{\rho\}$, where $\rho$ is the radical-pair density matrix. It can be seen from the master equation evolving $\rho$ (see caption of Fig. 2) that when calculating $d\rm Tr\{\rho\}/dt$, the Hamiltonian term that contains the magnetic field drops out. When the recombination rates are equal, as in the example of Fig. 2, it is $d\rm Tr\{\rho\}/dt=-k\rm Tr\{\rho\}$, and the sensitivity of $\rm Tr\{\rho\}$ to the magnetic field is identically zero. The authors in [127] observed a magnetic field effect because for ${\rm C^{\bullet+}PF^{\bullet-}}$ it is $k_{\rm S}\neq k_{\rm T}$, and the reaction terms for unequal recombination rates indirectly produce a magnetic-field dependence of $\rm Tr\{\rho\}$. Still, the magnetic sensitivity when measuring $\rm Tr\{\rho\}$ is quite lower than detecting e.g. the singlet product yield, as magnetoreceptive organisms are supposed to do physiologically. Additionally, molecular excitation is not as sharp as the laser probing of atoms, hence the measurement noise appears to be significant.

### v2

# Approaching the quantum limit of energy resolution in animal magnetoreception

@@ I Introduction
Numerous behavioural experiments have established the ability of many species to sense the geomagnetic field, yet the mechanisms behind biological magnetic sensing, and their working at the physical, chemical and physiological level, are still being debated. There is evidence that migratory birds [11, 12, 13, 14, 15], fruit flies [16, 17, 18, 19, 20, 21, 22, 23, 24], butterflies [25, 26, 27], turtles [28, 30, 29], sharks [31, 32, 33], eels [34, 35, 36, 37, 38], salmon [39, 40, 41], are among the many species having magnetic sensing capabilities, either at the level of a compass for acquiring directional information towards long-distance navigation, or in some form of a magnetic map for positioning, or both. Three mechanisms have so far been prevalent in the discussion of biological magnetoreception, the radical-pair [42, 43, 44, 45, 46], the magnetite [47, 48, 49, 50, 51], and the induction mechanism [52, 53, 54, 55], while a mechanism synthesizing the first two, the MagR protein complex, was relatively recently proposed [56].
@@
Figure 1: Energy resolution limit in magnetic sensing [85]. Magnetic sensitivity versus sensor’s linear dimension for tens of realizations of several magnetic sensing technologies, like atomic magnetometers, SQUIDs, diamond and Hall effects sensors, and more. The red solid line is defined by the energy resolution limit, that is, the bound $(\widetilde{\delta B})^{2}V/2\mu_{0}\geq\hbar$, where $\widetilde{\delta B}$ is the magnetic sensitivity (or equivalently magnetic noise) in units of ${\rm T/\sqrt{Hz}}$, $V$ is the volume of the sensor (the third power of the linear dimension), and $\mu_{0}$ vacuum’s magnetic permeability constant. If the measurement time is $\tau$, and the error in the magnetic field estimate obtained within the measurement time is $\delta B$, then $\widetilde{\delta B}=\delta B\sqrt{\tau}$. Figure reproduced from [85] with permission of the American Physical Society.
@@
Qualitatively, the closer to $\hbar$ the numerical value of the ER is for any given sensor, the more “quantum” can the sensor be characterized. This offers an alternative perspective in the study of quantum biology of magnetoreception [87, 88, 89, 90, 91, 92, 93, 94, 95, 96]. This discussion is not about a choice between “black” or “white”, since “quantumness” can now be quantified, and spans a continuum. Indeed, quantum information science has introduced quantifiers for quantum coherence [97] and entanglement [98], and some have been applied to quantifying “quantumness” of biological magnetoreception [99, 100]. The ER provides yet another quantifier, this time in units of $\hbar$. For example, if it is found that some biological magentic sensor has ER e.g. $10^{7}\hbar$, and thus is far from the quantum limit, this finding is still quite informative, since it implies room for $10^{7}$ improvement. Even if Nature did not have the chance to realize such an improvement, it could in principle be engineered with biomimetic sensors [101, 102, 103].
Quantitatively, the ERL provides a strong and subtle constraint of basic and measurable parameters ($\delta B$, $V$, $\tau$). This constraint is particularly helpful for understanding biological magnetoreceptors, which unavoidably face the complexities and uncertainties of biological systems, as opposed to precisely controlled quantum magnetometers in the laboratory. As will be outlined in the paper, some biological ERs cannot be attributed a specific value like the ERs of man-made sensors, but rather, a range of possible values, stemming from those uncertainties. Conversely, it will be seen that the ERL narrows down those uncertainties. For example, we here obtain a precise lower bound for the number of cryptochromes or MagR complexes that are required to obtain a given magnetic sensitivity within a given measurement time.
The outline of this paper is the following. In Sec. II we make some introductory comments about the energy resolution limit aimed at an interdisciplinary readership. We then apply the ERL to the radical-pair (Sec. III), the magnetite (Sec. IV), and the MagR mechanism (Sec. V). In Sec. VI we present the results in a global ERL map. We conclude in Sec. VII.
@@ II Introductory comments about the energy resolution limit
The left-hand-side defines the energy resolution of the sensor, and the right-hand side is Planck’s constant $\hbar\approx 10^{-34}~{}{\rm Js}$, which constant pervades quantum phenomena.
In this expression, $\delta B$ is the uncertainty in the sensor’s magnetic field estimate, $V$ the sensitive volume, and $\tau$ the measurement time. The constant $\mu_{0}=4\pi\times 10^{-7}~{}{\rm Tm/A}$ is the magnetic permeability of vacuum. The name “energy resolution” derives from the following. From classical electromagnetism we know that $B^{2}/2\mu_{0}$ is the magnetic energy density (unit ${\rm J/m^{3}}$) due to the presence of a magnetic field $B$ in space. For a homogeneous magnetic field, the total energy within some volume $V$ is then $B^{2}V/2\mu_{0}$. The ERL involves the variance $(\delta B)^{2}$ instead of $B^{2}$, therefore one can loosely think of it as the minimum magnetic energy measurable within the sensor volume during the time available for the measurement. The right-hand-side of (1) is approximately equal to $\hbar$, thus it is tempting to guess that some basic quantum principle underlies the ERL.
@@
In Fig. 1 we reproduce the results of [85], displaying the quantity $\widetilde{\delta B}=\delta B\sqrt{\tau}$ (having unit ${\rm T/\sqrt{Hz}}$) versus the linear dimension $L$ of the sensor. This is connected to the volume entering (1) by $V=L^{3}$. The spectral magnetic sensitivity $\widetilde{\delta B}$ describes how the measurement noise is distributed over the measurement bandwidth, given by $1/\tau$. The utility of $\widetilde{\delta B}$ is that we can produce a common figure for $\widetilde{\delta B}$ versus $L$ for a number of sensors having different measurement times $\tau$. It is seen in Fig. 1 that all sensors examined have $\widetilde{\delta B}\geq\sqrt{2\mu_{0}\hbar/L^{3}}$, which is another way to express the bound (1).
In the next three sections we analyze the ER of the radical-pair, the magnetite, and the MagR mechanism. We do not further discuss the induction mechanism, since the relevant ER stands at the level of $10^{26}\hbar$, rendering the mechanism purely classical, and thus far from any potential biomimetic advancement towards the quantum limit. In the beginning of each of the following three sections we briefly recapitulate the workings of the relevant mechanism. To establish the ER, we extract from the literature the magnetic sensitivity $\delta B$, the sensor volume $V$, and the measurement time $\tau$. At the end, we produce a similar map of the global results like the one in Fig. 1.
@@ III Radical-pair Mechanism
Within the perspective of the ERL we focus on estimating the magnetic field amplitude. This does not limit the utility of our results, since as stated in [6], “map and compass cues might not always be as separable as previously thought”. Information on the magnetic field change is reflected in the change in concentration of radical-pair reaction products. We remind the reader that such reactions involve a coherent spin motion, which is influenced by the magnetic field. This spin motion is driven by the intra-molecular magnetic fields produced by the hyperfine couplings of the two unpaired electrons of the donor and acceptor molecule forming the radical-pair with the magnetic nuclei of the respective molecule. The radical-pair is created by photo-excitation of the neutral precursor. The mechanism is depicted in Fig. 2a, which shows the photo-excitation step of the donor-acceptor molecular dyad, DA, the electron transfer creating the radical-pair state ${\rm D^{\bullet+}A^{\bullet-}}$ in the singlet state of the two unpaired electrons, the coherent spin motion ${\rm{}^{S}D^{\bullet+}A^{\bullet-}}\leftrightarrow{\rm{}^{T}D^{\bullet+}A^{\bullet-}}$ interchanging singlet with triplet radical-pair states, and the inverse electron transfer leading to the singlet (DA) and triplet (${\rm{}^{T}DA}$) reaction products.
@@
The signalling mechanism conveying the change in reaction yields to higher levels of neurophysiological processing is still unknown [126, 100]. Therefore, as magnetic-dependent observable one usually considers the relative proportion of singlet versus triplet reaction yields. The singlet (triplet) yield is found by integrating the singlet (triplet) probability of the radical-pair state times the probability for singlet (triplet) recombination, which for a time interval $dt$ is given by $k_{\rm S}dt$ ($k_{\rm T}dt$). The singlet (triplet) probability of the radical-pair state is theoretically calculated from the quantum state evolution of the radical-pair spin state (see caption of Fig. 2). The rates $k_{\rm S}$ and $k_{\rm T}$ are characteristic parameters of the specific radical-pair considered. In general they are different, however, in many simulations they are taken equal, $k_{S}=k_{T}\equiv k$. In this case, also considered herein, $\tau=1/k$ is the reaction time, which will also serve as the measurement time in our ER considerations.
@@
For completeness, in Fig. 2b we depict an example of the coherent spin motion ${\rm{}^{S}D^{\bullet+}A^{\bullet-}}\leftrightarrow{\rm{}^{T}D^{\bullet+}A^{\bullet-}}$ for an infinite radical-pair lifetime, i.e. $k=0$. We plot the singlet probability as a function of time for two different magnetic fields, to show how the field modulates the singlet-triplet mixing. In Fig. 2c we “turn on” the recombination channels, so in this case we observe the decay of the singlet probability, since radical-pairs recombine and their population tends to zero, while the population of the reaction products increases. This is shown in Fig. 2d for the singlet reaction yield, which is seen to convey information on the magnetic field. Indeed, for a change from $B=0$ to $B=0.5~{}{\rm G}$, the singlet reaction yield at the end of the reaction ($t\gg\tau$) changes by 10% (in absolute terms).
@@ III.1 Energy resolution of synthesized radical-pairs probed with laser spectroscopy
The authors in [127] synthesized a radical-pair model using a carotenoid-porphyrin-fullerene triad (CPF). The porphyrin serves as an intermediate cation in the electron-transfer process, whereas the actual radical-pair is ${\rm C^{\bullet+}PF^{\bullet-}}$, i.e. the carotenoid is the radical cation, and the fullerene is the radical anion. The authors observed a clear magnetic field effect in the radical-pair reaction at earth’s magnetic field, albeit at 113 K, i.e. much cooler than the physiological animal temperature. Nevertheless, these data are perfectly suitable to apply the ERL. The authors measured the transient absorption of the radical-pair state at $39~{}{\rm\mu T}$ and $49~{}{\rm\mu T}$. From the noise in their data (Fig. 2c of [127]) one can estimate the magnetic sensitivity to be $\delta B\approx 2.5~{}{\rm\mu T}$. The measurement time was $\tau\approx 1~{}{\rm\mu s}$, given by the duration of the magnetic-field effect. Finally, the sensing volume of the $100~{}{\rm\mu M}$ CPF solution that was probed by the laser measuring the transient absorption of ${\rm C^{\bullet+}PF^{\bullet-}}$ was $V\approx 2\times 10^{-8}~{}{\rm m^{3}}$ [128]. Thus in this case, it is ${\rm ER}=(\delta B)^{2}V\tau/2\mu_{0}=5\times 10^{14}\hbar$.
@@
The main reason for the large value of this ER is the large value of $\delta B$ deriving from the specific measurement scheme. It is known from quantum magnetometers working with atomic vapors [129] that the number of atoms participating in the measurement is crucial. The number density of CPF molecules, given their $100~{}{\rm\mu M}$ concentration, is $6\times 10^{16}~{}{\rm cm}^{-3}$, two orders of magnitude higher than the atom number density of typical optical pumping magnetometers [129], which deliver magnetic sensitivity at the level of 1 fT, with a corresponding ER at the level of $10\hbar$. Thus, $\delta B$ of CPF is not limited by the number of CPF molecules (or the photo-excited fraction thereof). Similarly, atomic magnetometers have about 3-4 orders of magnitude longer relaxation time, so again, $\tau$ is not the limiting factor here. Instead, it appears that the large value of the CPF energy resolution is dominated by the limited sensitivity of the specific measurement scheme.
@@
Indeed, the magnetic field effect is measured by detecting an optical absorption exciting the radical-pair state into higher electronically excited states, without any discrimination between singlet or triplet radical-pairs. That is, the magnetic field effect probes the total population of the radical-pair state, given by $\rm Tr\{\rho\}$, where $\rho$ is the radical-pair density matrix. It can be seen from the master equation evolving $\rho$ (see caption of Fig. 2) that when calculating $d\rm Tr\{\rho\}/dt$, the Hamiltonian term that contains the magnetic field drops out. When the recombination rates are equal, as in the example of Fig. 2, it is $d\rm Tr\{\rho\}/dt=-k\rm Tr\{\rho\}$, and the sensitivity of $\rm Tr\{\rho\}$ to the magnetic field is identically zero. The authors in [127] observed a magnetic field effect because for ${\rm C^{\bullet+}PF^{\bullet-}}$ it is $k_{\rm S}\neq k_{\rm T}$, and the reaction terms for unequal recombination rates indirectly produce a magnetic-field dependence of $\rm Tr\{\rho\}$. Still, the magnetic sensitivity when measuring $\rm Tr\{\rho\}$ is quite lower than when detecting e.g. the singlet product yield, as magnetoreceptive organisms are supposed to do physiologically. Additionally, molecular excitation is not as sharp as the laser probing of atoms, hence the measurement noise appears to be significant.